A piston-cylinder device containing carbon-dioxide gas undergoes an isobaric process from 103.4 kPa and 300 K to 366 K. Determine the work and heat transfer associated with this process in kJ/kg.
Answer #1


Q. CO2 is a diatomic gan. So, the degrees of freedom for the carbon dioxide is f = 3x3 - 2 = 7 and the adiabatic expansion coar and cp-cv =R 9 cp Cv = ... Cp - c =R & Cp (1-Ź)=R - Cp x(7-1)=R » Cp - (¥) R (10) Q = n(Cp-R) OT = n(NR - ROOT = n Rot (2-